OpenStudy (diamondboy):

What is the limit as x approaches 0 of (cotxsinx-1)/x

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~0}\frac{\cot(x)\sin(x)-1}{x}}\) like this?

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~0}\frac{\cot(x)\sin(x)-1}{x}}\) \(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~0}\frac{\cos(x)-1}{x}}\) then this is one of those theorems...

OpenStudy (diamondboy):

yea

OpenStudy (diamondboy):

yea I got to dat so what is the answer?

OpenStudy (solomonzelman):

The rule is, \(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~0}\frac{1-\cos(x)}{x}=0}\)

OpenStudy (diamondboy):

YES!!!!

OpenStudy (solomonzelman):

and cos(x)-1 is same as 1-cos(x) but times -1.

OpenStudy (diamondboy):

Then I got the answer right

OpenStudy (solomonzelman):

:) very ncie!

OpenStudy (diamondboy):

It came out in my test today

OpenStudy (solomonzelman):

ncie ? lol ... nice...!

OpenStudy (diamondboy):

Thank u

OpenStudy (solomonzelman):

yes, they are tricky at times.... just noticing the simplest approach, is the task:)

OpenStudy (solomonzelman):

yw

OpenStudy (diamondboy):

yep